Solve this equation.
sqrt(4z+1)=3+sqrt(4z-2)

Answer by user_dude2008(1861)

sqrt(4z+1)=3+sqrt(4z-2)
4z+1=(3+sqrt(4z-2))^2
4z+1=9+6*sqrt(4z-2)+4z-2
-6=6*sqrt(4z-2)
-1=sqrt(4z-2)
1=4z-2
4z=3
z=3/4

Check:
sqrt(4z+1)=sqrt(4(3/4)+1)=sqrt(3+1)=sqrt(4)=2
3+sqrt(4z-2)=3+sqrt(4(3/4)-2)=3+sqrt(3-2)=3+sqrt(1)=4

---> Contradiction


Answer: No Solutions